OpenStudy (anonymous):

write as a single logarithm a) log16+2log 3 b) log base 2 16 - 2logbase 2 3 c) 2+logbase 4 5

OpenStudy (anonymous):

A, 2log3=log3^2=log9 log16+log9=log16*9=144 B, 2log(base2) 3=log(bas2)9 log(base2)16-log(base2)9=log(base2)16/9 C, 2=log(base 4)4^2 log(base4)16+log(base4)5=log(base4)80

OpenStudy (anonymous):

thank you andras!